The FBI claims that about 80% of all property crimes, the criminals are never found and the case is never solved. In a randomly selected city, what is the probability that at least 1 crime will be solved when 6 crimes are reported?
I need help setting up the formulas

Answer by stanbon(75887)

The FBI claims that about 80% of all property crimes, the criminals are never found and the case is never solved.
In a randomly selected city, what is the probability that at least 1 crime will be solved when 6 crimes are reported?
I need help setting up the formulas
----------
Binomial Problem with n = 6 and p(never solved) = 0.8; p(solved) = 0.2
----
P(none solved)+P(at least one solved) = 1
P(at least one solved) = 1 - P(none solved)
---
= 1 - 6C0(0.2)^0*0.8^6 = 1 - 1*1*0.8^6 = 1 - 0.8^6 = 0.7389
==========================
Cheers,
Stan H.